# Large time behavior for the 3D Navier-Stokes with Navier boundary conditions

### v3

@@ Abstract.
We study the three-dimensional incompressible Navier-Stokes equations in a smooth bounded domain $\Omega$ with initial velocity $u_{0}$ square-integrable, divergence-free and tangent to $\partial\Omega$. We supplement the equations with the Navier friction boundary conditions $u\cdot\bm{n}=0$ and $[(2Su)\bm{n}+\alpha u]_{{\scriptstyle{\mathrm{tan}}}}=0$, where $\bm{n}$ is the unit exterior normal to $\partial\Omega$, $Su=(Du+(Du)^{t})/2$, $\alpha\in C^{0}(\partial\Omega)$ is the boundary friction coefficient and $[\cdot]_{{\scriptstyle{\mathrm{tan}}}}$ is the projection of its argument onto the tangent space of $\partial\Omega$. We prove global existence of a weak Leray-type solution to the resulting initial-boundary value problem and exponential decay in energy norm of these solutions when friction is positive. We also prove exponential decay if friction is non-negative and the domain is not a solid of revolution. In addition, in the frictionless case $\alpha=0$, we prove convergence of the solution to a steady rigid rotation, if the domain is a solid of revolution. We use the Galerkin method for existence, Poincaré-type inequalities, with suitable adaptations to account for the differential geometry of the boundary, and a global-in-time, integral Gronwall-type inequality.
@@ 1. Introduction
The Navier friction condition was first introduced in 1827 by C. Navier, see [Navier1827]. This boundary condition is often considered in the presence of rough boundaries, see [G-VM10, JM01] and references therein, or in flows with a turbulent layer next to the boundary, see [JLS04]. Physically, it expresses a rough balance between the rate of fluid slip and friction stress at the boundary. In addition, allowing for slip may be regarded as a relaxation of the no-slip boundary condition, which creates stiffness in computational modelling of high-Reynolds number flows, see [GS22a].
@@
In broad terms, we follow closely the corresponding results for $\alpha=\infty$. Existence is tackled by passing to the limit in a Galerkin approximation and the exponential decay is obtained by combining a Poincaré-type inequality with an integral Gronwall-type inequality. However, the Navier friction condition introduces a few technical hurdles that must be addressed. Specifically, we must account for the various equivalent weak formulations of the problem, the precise choice of basis for the Galerkin approximation, in a way that is consistent with the Navier condition, and the influence of the geometry of the boundary. In [CMR], Clopeau and collaborators addressed several of these issues in the two-dimensional case, proving global well-posedness of (1.1) and convergence of the vanishing viscosity limit. We make use of some of the ideas developed in [CMR], adapting and extending them as needed to the three-dimensional case.
In [BJ21], H. Baba and M. Jazar studied the long-time behavior of a weak solution to the Navier-Stokes system in a bounded 3D domain with Navier-type boundary conditions. In their work, the friction condition $[(2Su)\bm{n}+\alpha u]_{{\scriptstyle{\mathrm{tan}}}}=0$ is replaced by assuming vorticity is normal to the boundary. These types of boundary conditions are only equivalent in special cases, namely, if $\Omega$ is a half-space, and $\alpha=0$, or if $\Omega$ is a sphere of radius $r>0$ and $\alpha=2/r$. Moreover, their result only applies to initial data which satisfies their boundary condition and lies in the orthogonal complement of the kernel of the Stokes operator within that initial data space. Furthermore, they only obtain algebraic decay of the solution in energy norm.
@@ 3. The Stokes operator
$\leqslant C\|u\|_{H^{1}}\|v\|_{H^{1}}+\left(\|\lambda\|_{L^{\infty}(\Omega)}+\|\alpha\|_{L^{\infty}(\Omega)}\right)\|u\|_{L^{2}(\partial\Omega)}\|v\|_{L^{2}(\partial\Omega)}$
$\leqslant(C+\|\lambda\|_{L^{\infty}(\Omega)}+\|\alpha\|_{L^{\infty}(\Omega)})\|u\|_{H^{1}}\|v\|_{H^{1}},$
@@
We claim that there exists $C_{\beta}>0$ such that $\widetilde{B}:=B+C_{\beta}\mathbb{I}$, with $\mathbb{I}$ denoting the identity operator, is coercive with respect to the $H^{1}$-norm, i.e., there exists $c>0$ such that $\widetilde{B}(u,u)\geqslant c\|u\|_{H^{1}}^{2}$ for all $u\in H^{1}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)$. Indeed, this follows immediately from the estimate below:
@@
$\geqslant\|Du\|_{L^{2}}^{2}-(\|\alpha\|_{L^{\infty}(\Omega)}+\|\lambda\|_{L^{\infty}(\Omega)})\|u\|_{L^{2}(\partial\Omega)}^{2}$
$\geqslant\|Du\|_{L^{2}}^{2}-(\|\alpha\|_{L^{\infty}(\Omega)}+\|\lambda\|_{L^{\infty}(\Omega)})C\|u\|_{L^{2}(\Omega)}\|Du\|_{L^{2}(\Omega)}$
$\geqslant\frac{1}{2}\|Du\|_{L^{2}}^{2}-C\|u\|_{L^{2}(\Omega)}^{2}$
$\geqslant c\|u\|_{H^{1}}^{2}-C_{\beta}\|u\|_{L^{2}(\Omega)}^{2}.$ (3.5)
We have used, above, the trace inequality (3.3), the Poincaré inequality (3.4), and Young’s inequality.
@@
Let us introduce $\widetilde{\mathbb{A}}:=\mathbb{A}+C_{\beta}\mathbb{I}$, where $C_{\beta}$ is the constant from estimate (3). Consider $\text{Dom}(\widetilde{\mathbb{A}}):=\{u\in H^{1}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)\,|\,\widetilde{\mathbb{A}}u\in L^{2}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)\}$ and note that $\text{Dom}(\widetilde{\mathbb{A}})$ is compactly imbedded in $L^{2}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)$. Furthermore, it is easy to see that $\text{Dom}(\widetilde{\mathbb{A}})$ is dense in $L^{2}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)$ with respect to the $L^{2}$-norm.
The coerciveness of $\widetilde{B}$, together with the continuity and symmetry of $B$ and the compactness mentioned above, imply that the operator $\widetilde{\mathbb{A}}$ has an inverse which is a self-adjoint compact operator from $L^{2}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)$ to $\text{Dom}(\widetilde{\mathbb{A}})\subset H^{1}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)$. Therefore it follows from the spectral theory for compact self-adjoint operators that there exists an orthonormal basis of $L^{2}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)$ of eigenvectors of $\widetilde{\mathbb{A}}$, along with an increasing sequence of eigenvalues tending to +$\infty$+ ∞, see [BoyerFabrie2013, Section 6, Chapter II]. This basis is also orthogonal with respect to the ((,))(($\cdot$,$\cdot$))( ( ⋅ , ⋅ ) ) inner product on $H^{1}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)$, and it is a complete set, [BoyerFabrie2013, Corollary II.3.8 and Theorem IV.5.5]. Clearly, an eigenvector $v_{j}$ of $\widetilde{\mathbb{A}}$ is also an eigenvector of $\mathbb{A}$, albeit with eigenvalue $\lambda_{j}-C_{\beta}$. ∎
@@ 4. Existence of weak solution
$\frac{d}{dt}\|u^{M}\|_{L^{2}(\Omega)}^{2}\leqslant-2\nu\|Du^{M}\|_{L^{2}(\Omega)}^{2}+2\nu\left(\|\lambda\|_{L^{\infty}(\Omega)}+\|\alpha\|_{L^{\infty}(\Omega)}\right)\|u^{M}\|_{L^{2}(\partial\Omega)}^{2},$
@@
$\frac{d}{dt}\|u^{M}\|_{L^{2}(\Omega)}^{2}+\nu\|Du^{M}\|_{L^{2}(\Omega)}^{2}\leqslant C\nu\left(\|\lambda\|_{L^{\infty}(\Omega)}+\|\alpha\|_{L^{\infty}(\Omega)}\right)\|u^{M}\|_{L^{2}(\Omega)}^{2}.$ (4.2)
@@
$\frac{d}{dt}\|u^{M}\|_{L^{2}(\Omega)}^{2}\leqslant-2\nu\|Du^{M}\|_{L^{2}(\Omega)}^{2}+2\nu\int_{\partial\Omega}\left(\|\lambda\|_{L^{\infty}(\Omega)}-\alpha\right)|u^{M}|^{2}\,\mathrm{d}S.$ (4.9)
@@
$+2\nu\int_{s}^{r}\int_{\partial\Omega}\left(\|\lambda\|_{L^{\infty}(\Omega)}-\alpha\right)|u^{M}|^{2}\,\mathrm{d}S\,\mathrm{d}\tau+\|u^{M}(s,\cdot)\|_{L^{2}(\Omega)}^{2}.$ (4.10)
@@
$\leqslant 2\nu\int_{0}^{+\infty}|\phi_{\varepsilon}|^{2}(r)\int_{s}^{r}\int_{\partial\Omega}\left(\|\lambda\|_{L^{\infty}(\Omega)}-\alpha\right)|u^{M}|^{2}\,\mathrm{d}S\,\mathrm{d}\tau\,\mathrm{d}r+\|u^{M}(s,\cdot)\|_{L^{2}(\Omega)}^{2}.$ (4.11)
@@
$\leqslant 2\nu\int_{0}^{+\infty}|\phi_{\varepsilon}|^{2}(r)\int_{s}^{r}\int_{\partial\Omega}\left(\|\lambda\|_{L^{\infty}(\Omega)}-\alpha\right)|u|^{2}\,\mathrm{d}S\,\mathrm{d}\tau\,\mathrm{d}r$
@@
$\leqslant 2\nu\int_{s}^{t}\int_{\partial\Omega}\left(\|\lambda\|_{L^{\infty}(\Omega)}-\alpha\right)|u|^{2}\,\mathrm{d}S\,\mathrm{d}\tau+\liminf_{M\to+\infty}\|u^{M}(s,\cdot)\|_{L^{2}(\Omega)}^{2}.$ (4.12)
@@ 5. Symmetric Poincaré inequality
In Theorem 4.4 we obtained the existence of a weak solution to (1.1) which satisfies two different energy inequalities. In the remainder of this article, we will be showing that, under appropriate hypotheses on the friction coefficient $\alpha$, we can prove exponential decay of the weak solution in energy norm. A crucial tool to prove such results are Poincaré-type inequalities. In this section we obtain such an inequality in terms of the symmetric part of the Jacobian of $u$; this is related to Korn’s inequality in $H^{1}_{0}$, see [BoyerFabrie2013, Remark IV.7.3, (IV.87)].
@@ 7. Exponential decay – Part 2
The following result is an immediate consequence of Lemmas 7.1 and 7.2.

### v4

@@ Abstract.
We study the three-dimensional incompressible Navier-Stokes equations in a smooth bounded domain $\Omega$ with initial velocity $u_{0}$ square-integrable, divergence-free and tangent to $\partial\Omega$. We supplement the equations with the Navier friction boundary conditions $u\cdot\bm{n}=0$ and $[(2Su)\bm{n}+\alpha u]_{{\scriptstyle{\mathrm{tan}}}}=0$, where $\bm{n}$ is the unit exterior normal to $\partial\Omega$, $Su=(Du+(Du)^{t})/2$, $\alpha\in C^{0}(\partial\Omega)$ is the boundary friction coefficient and $[\cdot]_{{\scriptstyle{\mathrm{tan}}}}$ is the projection of its argument onto the tangent space of $\partial\Omega$. We prove global existence of a weak Leray-type solution to the resulting initial-boundary value problem and exponential decay in energy norm of these solutions when friction is positive. We also prove exponential decay if friction is non-negative and the domain is not a solid of revolution. These two results are well known in the case of Dirichlet boundary condition, but, even if they have been implicitly used for the Navier boundary conditions, the comprehensive analysis is not available in the literature. After carefully studying the Stokes semigroup for such a boundary condition, we use the Galerkin method for existence, Poincaré-type inequalities, with suitable adaptations to account for the differential geometry of the boundary, and a novel integral Gronwall-type inequality. In addition, in the frictionless case $\alpha=0$, we prove convergence of the solution to a steady rigid rotation, if the domain is a solid of revolution.
@@ 1. Introduction
The Navier friction condition was first introduced in 1827 by C. Navier, see [Navier1827]. This boundary condition is often considered in the presence of rough boundaries, see [G-VM10, JM01] and references therein, or in flows with a turbulent layer next to the boundary, see [JLS04]. Physically, it expresses a rough balance between the rate of fluid slip and friction stress at the boundary. In addition, allowing for slip may be regarded as a relaxation of the no-slip boundary condition, which creates stiffness in computational modeling of high-Reynolds number flows, see [GS22a].
The standard no-slip boundary condition $u=0$ corresponds to formally setting $\alpha=\infty$ above. In this case, global-in-time existence of a weak Leray-Hopf solution, for initial data $u_{0}$ divergence-free, tangent to the boundary and square-integrable, is due to J. Leray, see [L34a]. Additionally, the exponential decay of the solution in $L^{2}$ follows from the energy inequality, by using the Poincaré inequality.
In the case $0<\alpha<\infty$ the literature contains no complete proof, aside from some special cases, of existence of weak (Leray-Hopf) solutions or discussion of large time behavior. Global existence of weak solutions was proved in the case $\alpha=0$ in [FG2022DCDS, Watanabe2003]. A sketch of a proof for $0\leqslant\alpha<\infty$ was given in [IS2011]. The rigorous analysis for global existence was presented for several boundaries conditions in [BoyerFabrie2013] but the Navier boundary condition is not covered. In [AACG2021, AEG2021] local well-posedness of strong solutions was established using semigroup methods and mild solutions, i.e. in which the equations are written in integral form, under appropriate regularity hypotheses on $\alpha$ and assuming the initial velocity is in $L^{r}$, $r\geqslant 3$. Global existence, under these conditions, was obtained for small data. For two-dimensional domains, Clopeau, Mikelić and Robert proved global well-posedness of weak solutions and convergence of the vanishing viscosity limit in [CMR]. Their analysis is based on the vorticity $\omega:=\operatorname{curl}u$ and the stream function $\psi:=\Delta^{-1}_{0}\omega$ (solution of the Laplace problem with the Dirichlet boundary condition) and hence applied only to 2D, simply connected, smooth, bounded domains. Taking into account harmonic vector fields and circulations, the first author extended their result to domains with holes in [KNavier]. The goal here is to present a complete and self-contained analysis for 3D bounded domains.
@@
In broad terms, we follow the general strategy developed for similar results when $\alpha=\infty$. Existence is tackled by passing to the limit in a Galerkin approximation and the exponential decay is obtained by combining a Poincaré-type inequality with an integral Gronwall-type inequality. However, the Navier friction condition introduces a few technical hurdles that must be addressed. Specifically, we must account for the various equivalent weak formulations of the problem, the precise choice of basis for the Galerkin approximation, in a way that is consistent with the Navier condition, and the influence of the geometry of the boundary.
In [BJ21], H. Baba and M. Jazar studied the long-time behavior of a weak solution to the Navier-Stokes system in a bounded 3D domain with Navier-type boundary conditions. In their work, the friction condition $[(2Su)\bm{n}+\alpha u]_{{\scriptstyle{\mathrm{tan}}}}=0$ is replaced by assuming vorticity is normal to the boundary. These types of boundary conditions are only equivalent in special cases, namely, if $\Omega$ is a half-space, and $\alpha=0$, or if $\Omega$ is a sphere of radius $r>0$ and $\alpha=2/r$. Moreover, their result only applies to initial data which satisfies their boundary condition and lies in the orthogonal complement of the kernel of the Stokes operator within that initial data space. Furthermore, they only obtain algebraic decay of the solution in energy norm. Recently a study on the long-time asymptotics of strong solutions with $\alpha=0$ has become available, see [BFP24], which rediscovers some of the details in the present work.
Additional related work on Navier-Stokes and Stokes with Navier-boundary conditions include [SSc1973, Berselli2010, BdVeiga2004] and references therein.
@@ 3. The Stokes operator
$\leqslant C\|u\|_{H^{1}}\|v\|_{H^{1}}+\left(\|\lambda\|_{L^{\infty}(\partial\Omega)}+\|\alpha\|_{L^{\infty}(\partial\Omega)}\right)\|u\|_{L^{2}(\partial\Omega)}\|v\|_{L^{2}(\partial\Omega)}$
$\leqslant(C+\|\lambda\|_{L^{\infty}(\partial\Omega)}+\|\alpha\|_{L^{\infty}(\partial\Omega)})\|u\|_{H^{1}}\|v\|_{H^{1}},$
@@
We claim that there exists $\beta>0$ such that $\widetilde{B}:=B+\beta\mathbb{I}$, with $\mathbb{I}$ denoting the identity operator, is coercive with respect to the $H^{1}$-norm, i.e., there exists $K>0$ such that $\widetilde{B}(u,u)\geqslant K\|u\|_{H^{1}}^{2}$ for all $u\in H^{1}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)$. Indeed, this follows immediately from the estimate below:
@@
$\geqslant\|Du\|_{L^{2}}^{2}-(\|\alpha\|_{L^{\infty}(\partial\Omega)}+\|\lambda\|_{L^{\infty}(\partial\Omega)})\|u\|_{L^{2}(\partial\Omega)}^{2}$
$\geqslant\|Du\|_{L^{2}}^{2}-(\|\alpha\|_{L^{\infty}(\partial\Omega)}+\|\lambda\|_{L^{\infty}(\partial\Omega)})C\|u\|_{L^{2}(\Omega)}\|Du\|_{L^{2}(\Omega)}$
$\geqslant\frac{1}{2}\|Du\|_{L^{2}}^{2}-\beta\|u\|_{L^{2}(\Omega)}^{2}$
$\geqslant K\|u\|_{H^{1}}^{2}-\beta\|u\|_{L^{2}(\Omega)}^{2}.$ (3.5)
Above, in the second inequality we used the trace inequality (3.3). In the third inequality we used Young’s inequality with
$\beta=\frac{(\|\alpha\|_{L^{\infty}(\partial\Omega)}+\|\lambda\|_{L^{\infty}(\partial\Omega)})^{2}C^{2}}{2}.$
Finally, in the fourth inequality we used the Poincaré inequality (3.4), where $K=(2c^{2})^{-1}$.
@@
Let us introduce $\widetilde{\mathbb{A}}:=\mathbb{A}+\beta\mathbb{I}$, where $\beta$ is the constant from estimate (3). Consider $\text{Dom}(\widetilde{\mathbb{A}}):=\{u\in H^{1}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)\,|\,\widetilde{\mathbb{A}}u\in L^{2}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)\}$ and note that $\text{Dom}(\widetilde{\mathbb{A}})$ is compactly imbedded in $L^{2}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)$. Furthermore, it is easy to see that $\text{Dom}(\widetilde{\mathbb{A}})$ is dense in $L^{2}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)$ with respect to the $L^{2}$-norm.
The coerciveness of $\widetilde{B}$, together with the continuity and symmetry of $B$ and the compactness mentioned above, imply that the operator $\widetilde{\mathbb{A}}$ has an inverse which is a self-adjoint compact operator from $L^{2}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)$ to $\text{Dom}(\widetilde{\mathbb{A}})\subset H^{1}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)$. Therefore it follows from the spectral theory for compact self-adjoint operators that there exists an orthonormal basis of $L^{2}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)$ of eigenvectors of $\widetilde{\mathbb{A}}$, along with an increasing sequence of eigenvalues tending to +$\infty$+ ∞, see [BoyerFabrie2013, Section 6, Chapter II]. This basis is also orthogonal with respect to the ((,))(($\cdot$,$\cdot$))( ( ⋅ , ⋅ ) ) inner product on $H^{1}_{\sigma,{\scriptstyle{\mathrm{tan}}}}(\Omega)$, and it is a complete set, [BoyerFabrie2013, Corollary II.3.8 and Theorem IV.5.5]. Clearly, an eigenvector $v_{j}$ of $\widetilde{\mathbb{A}}$ is also an eigenvector of $\mathbb{A}$, albeit with eigenvalue $\lambda_{j}-\beta$. ∎
@@ 4. Existence of weak solution
$\frac{d}{dt}\|u^{M}\|_{L^{2}(\Omega)}^{2}\leqslant-2\nu\|Du^{M}\|_{L^{2}(\Omega)}^{2}+2\nu\left(\|\lambda\|_{L^{\infty}(\partial\Omega)}+\|\alpha\|_{L^{\infty}(\partial\Omega)}\right)\|u^{M}\|_{L^{2}(\partial\Omega)}^{2},$
@@
$\frac{d}{dt}\|u^{M}\|_{L^{2}(\Omega)}^{2}+\nu\|Du^{M}\|_{L^{2}(\Omega)}^{2}\leqslant C\nu\left(\|\lambda\|_{L^{\infty}(\partial\Omega)}+\|\alpha\|_{L^{\infty}(\partial\Omega)}\right)\|u^{M}\|_{L^{2}(\Omega)}^{2}.$ (4.2)
@@
$\frac{d}{dt}\|u^{M}\|_{L^{2}(\Omega)}^{2}\leqslant-2\nu\|Du^{M}\|_{L^{2}(\Omega)}^{2}+2\nu\int_{\partial\Omega}\left(\|\lambda\|_{L^{\infty}(\partial\Omega)}-\alpha\right)|u^{M}|^{2}\,\mathrm{d}S.$ (4.9)
@@
$+2\nu\int_{s}^{r}\int_{\partial\Omega}\left(\|\lambda\|_{L^{\infty}(\partial\Omega)}-\alpha\right)|u^{M}|^{2}\,\mathrm{d}S\,\mathrm{d}\tau+\|u^{M}(s,\cdot)\|_{L^{2}(\Omega)}^{2}.$ (4.10)
@@
$\leqslant 2\nu\int_{0}^{+\infty}|\phi_{\varepsilon}|^{2}(r)\int_{s}^{r}\int_{\partial\Omega}\left(\|\lambda\|_{L^{\infty}(\partial\Omega)}-\alpha\right)|u^{M}|^{2}\,\mathrm{d}S\,\mathrm{d}\tau\,\mathrm{d}r+\|u^{M}(s,\cdot)\|_{L^{2}(\Omega)}^{2}.$ (4.11)
@@
$\leqslant 2\nu\int_{0}^{+\infty}|\phi_{\varepsilon}|^{2}(r)\int_{s}^{r}\int_{\partial\Omega}\left(\|\lambda\|_{L^{\infty}(\partial\Omega)}-\alpha\right)|u|^{2}\,\mathrm{d}S\,\mathrm{d}\tau\,\mathrm{d}r$
@@
$\leqslant 2\nu\int_{s}^{t}\int_{\partial\Omega}\left(\|\lambda\|_{L^{\infty}(\partial\Omega)}-\alpha\right)|u|^{2}\,\mathrm{d}S\,\mathrm{d}\tau+\liminf_{M\to+\infty}\|u^{M}(s,\cdot)\|_{L^{2}(\Omega)}^{2}.$ (4.12)
@@ 5. Symmetric Poincaré inequality
In Theorem 4.4 we obtained the existence of a weak solution to (1.1) which satisfies two different energy inequalities. In the remainder of this article, we will be showing that, under appropriate hypotheses on the friction coefficient $\alpha$, we can prove exponential decay of the weak solution in energy norm. A crucial tool to prove such results are Poincaré-type inequalities. In this section we state and prove such an inequality in terms of the symmetric part of the Jacobian of $u$; this is related to Korn’s inequality in $H^{1}_{0}$, see [BoyerFabrie2013, Remark IV.7.3, (IV.87)]. This result is already known, see [AR2014, Lemma 3.3]; we include a proof for the sake of completeness.
@@ 7. Exponential decay – Part 2
The following result is an immediate consequence of Lemmas 7.1 and 7.2. This result may also be found in [FG2022Annali, Theorem 1].
@@
The result above was originally obtained in [Watanabe2003, Theorem 6.2], albeit with a different proof. In particular, the aforementioned proof did not involve the result established in Proposition 7.4, namely, conservation of $\mathrm{Proj}_{\mathrm{Ker}\,S}u(t)$.